.An investor is bearish on a stock whose market price is K100 per share. He instructs his broker to sell short 1000 shares who borrows the 1000 either from another customer’s account or from another broker. Suppose the broker has a 50% margin requirement on short sales, how would the investor’s account look like? 


1
Expert's answer
2021-04-01T19:04:54-0400

"P=100"

"q=1000"

"\frac{50}{100} \times 100 \times 1000 = 50000"